Can i get a little help with this please
E10-3 (Acquisition Costs of Trucks) Kelly Clarkson Corporation operates a retail computer store. To improve delivery services to customers, the company purchases four new trucks on April 1, 2014. The terms of acquisition for each truck are described below 1. Truck #1 has a list price of-$15,000 and is acquired for a cash payment of $13 900 2 Truck #2 has a list price of $16 cash and a zero-interest bearing note with a face amount of April 1, 2015, Clarkson would normally have to pay interest at a rate of borrowing, and the dealership has an incremental borrowing rate of 3. Truck #3 has a list price of S16000 it is acquired inexcha Clarkson carries in inventory The computer system costs $12 000 and normally sold by Clarkson for $15,200 Clarkson uses perpetual inventory system 4. Truck #4 has a list price of l s 14 000 shares of common stock in Clarkson Corporation. The stock has a par value per share of .000 and is acquired for a down payment of $2 14000 The note is due 0% for such a nge for a computer system that It is acquired inexchange for 1.000 8% 、 ) $10 and a market value of per share Prepare the appropriate journal entries for the foregoing transactions for Clarkson Corporation. Amount Account Title Amount Amount Amount Account Title Account Title Account Title Amount Hint: Use the Excel Present Value formula. Cells are formatted to show to 2 decimal places. Account Title Account Title Amount Amount Account Title Account Title Amount Amou Account title Amount Account title Amount
Answer #1
No. Account Titles and Explanation Debit Credit
1 Truck $ 13,900.00
        Cash $    13,900.00
2 Trucks = PV(10%,1,0,-14000) + 2000 $ 14,727.27
Discount on Notes Payable $   1,272.73
                      Cash $      2,000.00
                     Notes Payable $    14,000.00
3 Trucks $ 15,200.00
Cost of Good Sold $ 12,000.00
                 Inventory $    12,000.00
                 Sales $    15,200.00
4 Truck $ 13,000.00
           Common Stock $    10,000.00
            Paid in Capital $      3,000.00
No. Account Titles and Explanation Debit Credit
1 Truck 13900
        Cash =C10
2 Trucks = PV(10%,1,0,-14000) + 2000 =(PV(10%,1,0,-14000))+2000
Discount on Notes Payable =D16+D15-C13
                      Cash 2000
                     Notes Payable 14000
3 Trucks 15200
Cost of Good Sold 12000
                 Inventory =C19
                 Sales =C18
4 Truck =1000*13
           Common Stock =1000*10
            Paid in Capital =C23-D24